OpenStudy (kanwal32):

If A and B are two acute angles satisfying the equation 3sin^2A + 2sin^2B=1 and 3sin2A-2sin2B=0 then A+2B

OpenStudy (goformit100):

Welcome to OpenStudy Pajji

OpenStudy (goformit100):

Sir Please Show your working.

OpenStudy (kanwal32):

I'll attach a file wait for once I have converted sorry I can't attch the file but I will till you the steps till where I reached 3sin^2A=1- 2sin^2B (1-2sin^2B=cos2b 3sin2A=2sin2b cancelling sin on both sides(I think this method is wrong) then getting 3A=2B then equating it in the above equation

OpenStudy (kanwal32):

Now what to do???????????? big question

OpenStudy (kanwal32):

Sir pls help

OpenStudy (kanwal32):

three anwers are given in the textbook for this question 14 ,15 and 17

OpenStudy (goformit100):

it's ok. Now solve and Show this question too in the same way :- 3sin2A-2sin2B=0

OpenStudy (kanwal32):

can I cancel sin on both sides?

OpenStudy (kanwal32):

or should I use double angle property

OpenStudy (goformit100):

not cancel only you can remove sine both sides

OpenStudy (goformit100):

@ParthKohli HELP BRO I got stuck in this question

OpenStudy (goformit100):

@patrickthestar

OpenStudy (kanwal32):

pls help @ParthKohli sir

OpenStudy (anonymous):

with?

OpenStudy (kanwal32):

I am not able to solve this problem I am trying it for last 1 hr pls any1 help because day after tomorrow is my exam

OpenStudy (kanwal32):

@patrickthestar have u to any step pls hare I will try to solve it from there

OpenStudy (kanwal32):

hare=share

OpenStudy (anonymous):

Hey, I think I found the value for 2a+2b..but that doesn't solve your purpose..does it?

OpenStudy (kanwal32):

pls share it

OpenStudy (anonymous):

It's around 110 degree..what is the answer?

OpenStudy (shadowlegendx):

@kanwal32 not many people are on who can answer these type of questions. I'm sorry but you may have to wait till tomorrow to have this question answered. @Hero @Luigi0210 Help if you can guys :)

OpenStudy (kanwal32):

90 is degree

OpenStudy (anonymous):

Let me check that again!

OpenStudy (anonymous):

I am getting 2a+2b=107.45 degree!

OpenStudy (luigi0210):

@dan815 will be better help :)

OpenStudy (shadowlegendx):

@dan815 @dan815 @dan815 We call upon thy booty @dan815 @dan815 @dan815 @dan815

OpenStudy (kanwal32):

@dan815 pls help

OpenStudy (dan815):

hey okay

OpenStudy (kanwal32):

pls save my life @dan815

OpenStudy (kanwal32):

thnx

OpenStudy (dan815):

sorry i believe i mis read your question the square is only for the first equations right

OpenStudy (kanwal32):

yes

OpenStudy (kanwal32):

can we conver 1-2sin^2(b) =cos2b

OpenStudy (dan815):

3sin^2A + 2sin^2B=1 <--- use this identitiy sin^2(A)=1/2(1-cos(2a)) 3sin2A-2sin2B=0

OpenStudy (dan815):

Yes and then derive the first line and then you will have it in terms of sin

OpenStudy (dan815):

Then you can use a linear combination to eliminate

OpenStudy (dan815):

3/2-3/2cos(2a)+1-cos2b=1 deriving this gives 3sin2a+2sin2b=0 other eqn 3sin2A-2sin2B=0 How we can add them together and eliminate 2sin2b and solve for A, remember A<pi/2 rad or <90 deg

OpenStudy (dan815):

Any questions?

OpenStudy (kanwal32):

no thnx